# Elastic Constants and Bending Rigidities from Long-Wavelength Perturbation Expansions

### v1

@@ IV.6 Elastic tensor and bending rigidity in semiconducting monolayers
To validate the generality of our bending rigidity approach, we also tested semiconducting monolayers with long-range fields that couple with lattice vibrations. We start with h-BN but remark that the the calculation of quadrupole tensor in DFPT is not implemented in Quantum ESPRESSO [104, 105] and the the charge density response is limited to in-plane atomic displacements [118]. Therefore, we use the quadrupole tensors for h-BN, MoS2 and InSe computed by some of us [98] using the DFPT implementation in the ABINIT package [157, 97, 158, 159] and reported in Table S1 of the SM [132]. To guarantee result consistency, we therefore use the same computational parameters as Ref. [98]. As a side remark, we find that the use of 2D Fourier interpolation with the long-range presented in Sec. III.4 combined with the rotational invariance and equilibrium conditions [90] gives small imaginary or linear ZA modes in the long-wavelength limit due to out-of-plane dynamical dipoles, the second term in square parentheses in Eq. (79). We find that these out-of-plane dipoles have negligible contributions to both the phonon dispersions and elastic properties in the studied materials and leave this limitation for further studies. We give additional details on the rotational invariance and equilibrium conditions in App. A.
@@
Figure 6: (a) Phonon dispersion and (b) independent components of the elastic and bending rigidity tensors in h-BN as a function of supercell size. The multipolar interactions up to the quadrupole terms are considered for the long-range effect in the phonon dispersion of h-BN, and the inset of panel (a) shows the zoom-in acoustic branches around the $\Gamma$ point along the K and M directions. In the panel (b), the short-circuit elastic constants and bending rigidities of h-BN are calculated at the levels of removing DD and DD+DQ+QQ interactions; the dotted lines represent the theoretical elastic constants $C_{11}$/$C_{12}$ and bending rigidity $D_{11}$ taken from Ref. [148] and Ref. [84], respectively.
@@
Figure 7: Phonon dispersions and independent components of the elastic and bending rigidity tensors in InSe as a function of supercell size. The phonon spectra in the panels (a) and (b) are obtained based on a Fourier interpolation with the DD and DD+DQ+QQ interactions taken into account, respectively, and their insets show the zoom-in acoustic branches around the $\Gamma$ point along the K and M directions. In the panel (c), the short-circuit elastic constants and bending rigidities of InSe are also calculated at the levels of removing DD and DD+DQ+QQ interactions; the dotted lines represent the theoretical elastic constants $C_{11}$ and $C_{12}$ taken from Ref. [155].
@@
When it comes to the bending rigidities of h-BN shown in Fig. 6(b), the convergence is faster than for elastic constants and reached for a 6$\times$6$\times$1 supercell which is connected to fast convergence of the long-wavelength dispersion of ZA modes as shown in the inset of Fig. 6(a). However, we notice a small drift at higher supercell sizes for the bending rigidities because in monolayers the necessary condition of an $n$-order moment of IFCs to converge is that the IFCs themselves decay faster than $1/d^{n+2}$, as the integral in Eq. (73) is performed over the in-plane area of a supercell. This implies that the decay of short-range IFCs in real space must be faster than $1/d^{6}$ in order to converge the fourth-order moment of IFCs in 2D materials, which is equivalent to the requirement that all the multipolar interactions at $\mathcal{O}(q^{4})$ should be be removed; these include the first-order DD, the second-order DQ, the third-order QQ and DO, and the fourth-order QO and DH interactions, where H denotes the hexadecapoles. We leave the calculations of multipoles moments higher than quadrupoles in 2D for further studies but notice that our results range within 1.01–1.06 eV, which are still in good agreement with the existing theoretical results, 1.02 eV [84] and 0.95 eV [149]. Also, interestingly, the effects of DQ and QQ interactions on bending rigidities are negligible, as shown in Fig. 6(b). Besides, we predict for the first time $D_{12}=0.28$ eV and a Gaussian modulus of -0.79 eV in h-BN. Lastly, unlike graphene, the lattice relaxation does contribute to the bending rigidities of h-BN, that is, -0.11 eV for both $D_{11}$ and $D_{12}$, which is negative and thus lowers the energy of the system upon bending.
@@
We continue to study the case of 2D MoS2 monolayer, whose detailed results are shown in Fig. S4 of the SM [132]. We find that convergence for that material is easier because the macroscopic long-range effects are weak, which we attribute to the relatively large in-plane dielectric permittivity but accompanied with small effective charges. Based on the IFCs of a 12$\times$12$\times$1 supercell, we obtain the $C_{11}$ and $C_{12}$ of monolayer MoS2 as 131.88 and 32.28 N/m, respectively, while the previous DFT values are 132.3 and 32.8 N/m [152]. The detailed mechanical and elastic properties of 2D MoS2 are summarized in Table 4, in excellent agreement with experiment. In addition, the converged principal bending rigidity $D_{11}$ based on a 12$\times$12$\times$1 supercell is 9.06 eV, consistent with a value of 9.01 eV [83] and 9.10 eV [81], which both relies on the Helfrich Hamiltonian method. Notably, we find that the lattice-mediated contribution due to relaxation effects significantly impacts the bending rigidities of MoS2, which is a crucial difference from the behavior observed in graphene and h-BN. Without the lattice-mediated contribution, both $D_{11}$ and $D_{12}$ become negative (i.e. -1.53 and -0.51 eV), which indicates that such material is not stable under a bending perturbation due to the failure of the clamped-ion approximation. It further suggests that the motions of bending waves in MoS2 can be only stabilized by the renormalization effect of the lattice, a phenomenon that has not been discussed before. Finally, we here report for the first time the values of $D_{12}$ and Gaussian modulus $D^{\mathrm{G}}$ of MoS2 are 4.35 and -4.71 eV, respectively, awaiting the future experimental verifications.
@@
Our last example for demonstrating the accuracy of the proposed method is monolayer InSe, which serves as a prototype for the 2D family of III–VI metal chalcogenides. The calculated phonon dispersions with the increasing supercell size are shown in Figs. 7(a) and (b), where the long-range Coulomb electrostatics is treated at the levels of DD and DD+DQ+QQ interactions, respectively. In the case of DD interaction only, the slope of the linear TA modes around the $\Gamma$ point changes slowly as the supercell size increases. As can be seen from Fig. 7(b) and the inset therein, when the DD+DQ+QQ interpolation scheme is used, the resulting phonon spectra can be readily converged using only a small 6$\times$6$\times$1 supercell. In particular, the long-wavelength behaviors of all three acoustic modes remain constant with the increasing supercell size from 4$\times$4$\times$1 to 14$\times$14$\times$1. For the short-circuit elastic constants, as illustrated in Fig. 7(c), the values of $C_{11}$ and $C_{12}$ are 49.60 and 20.43 N/m, respectively. While $C_{11}$ agrees well with the reference result of 49.38 N/m from the previous DFT calculations [155], $C_{12}$ shows a large error around 48.91% with respect to the reference value of 13.72 N/m. Such an inconsistency can be remedied by further taking into account the contributions from DQ and QQ interactions, as shown by the red dotted lines in Fig. 7(c). When the mutipolar interactions up to the quadrupole terms are considered, values of 49.27 and 13.38 N/m are obtained, respectively, in excellent agreement with the previous reported values. Also, if the clamped-ion approximation is adopted, the respective $C_{11}$ and $C_{12}$ are 73.38 and 23.39 N/m. The detailed results of elastic constants and moduli for 2D InSe are summarized in Table 4.
@@
Similarly to the cases of h-BN and MoS2, the bending rigidities of InSe in Fig. 7(c) suffer from the issue of divergence when using a large supercell, due to the existence of high-order mutipolar interactions beyond the quadrupole terms. Although the removal of DQ and QQ interactions only results in a small positive shift in the $D_{12}$ component compared to the case with only DD interactions subtracted, the bending rigidities are still diverging, especially from the supercell size of 12$\times$12$\times$1. We note that the impact of DD term itself is also relatively small, see Fig. S5 of the SM [132], which suggests that octupole tensor will be important in this material. However, owing to the technical difficulty of computing higher-order multipolar interactions, we will leave this point for the future investigation. Using a moderate 10$\times$10 supercell, our approximated principal bending rigidity $D_{11}$, Gaussian modulus $D^{\mathrm{G}}$ and the $D_{12}$ component of InSe monolayer are 22.60, -14.25 and 8.34 eV, respectively. By comparison, the principal bending rigidity of InSe from the Helfrich Hamiltonian method within the cyclic DFT calculations is 17.5 eV [83]. Such discrepancy could be attributed to a relatively large curvature used in their calculations, which can induce a in-plane stain field and the related atomic conformation relaxation as a drawback of the Helfrich Hamiltonian method. Furthermore, our simulation also reveals a significant renormalization effect from the lattice relaxation on bending rigidities of InSe, where the clamped-ion $D_{11}$ and $D_{12}$ are found to be 0.32 and 0.11 eV. Interestingly, for 2D materials containing only a single layer such as h-BN the lattice relaxation induced by a bending deformation tends to lower the bending rigidity, while it can significantly stiffen the bending modulus of multilayer 2D materials. This finding will provide practical guidelines for designing future flexible electronics.
@@ V Conclusions
In this work, we propose a novel method to simultaneously compute the elastic constants and bending rigidities of bulk and LD materials. Compared to existing methods, the approach is based on the long-wavelength perturbation theory and uses only the real-space IFCs, dynamical charges and dielectric tensors as input. This provides the direct benefit to work for any dimensions unlike the finite-difference method based on the strain-stress/energy relation where the elastic constants involving the out-of-plane Cartesian component should be removed in the constitutive equations for 2D materials. Second, the electromechanical coupling is treated on the same footing through the calculated multipole and dielectric tensors. One can readily obtain the open-circuit elastic constants via the piezoelectric tensor calculated directly from the quadrupole tensor, with the lattice-mediated contribution fully included. Third, the bending rigidities of 2D materials can be easily obtained based on a higher-order long-wavelength perturbation expansion in the same manner. Notably, the calculated bending rigidity is inherently tensorial, which further allows us to explore the elastic anisotropy of 2D plates, and currently this is the only way to obtain the complete fourth-rank bending rigidity tensor, to the best of our knowledge. Fourth, temperature-dependent elastic constants and bending rigidities beyond the quasiharmonic approximation can be readily achieved via the self-consistent harmonic approximation [160] by taking into account the full quantum anharmonicity, as our approach takes the real-space IFCs as the input. Lastly, if the phonon calculations are already performed, the calculations of elastic constants and bending rigidities by our approach can be performed at virtually no additional computational cost. This will enable the automated high-throughput calculations of elastic properties directly through the existing phonon workflow that has been already implemented in most of high-throughput workflow managers [161, 162].
@@
In summary, we present an efficient and accurate approach for calculating both the elastic and bending rigidity tensors of crystalline solids, which is rigorously formulated based on the secular equation of lattice dynamics and the long-wavelength perturbation theory. In particular, we have provided a practical recipe in the framework of modern electronic structure calculations for implementing the Huang formula of elastic stiffness tensor that dates back to the atomic theory of elasticity introduced in 1949. Besides, we extend the Huang’s theory to describe also the equations of motion of bending wave in 2D materials and obtain the corresponding bending rigidity tensor, through the fourth-order long-wavelength expansion. Significantly, in semiconductors and insulators, the induced macroscopic electric fields couple with the lattice vibrations in the long-wavelength limit, which must be treated separately in order to recover the elastic constants and bending rigidities at the correct electrical boundary conditions. To obtain the short-circuit one, it is found that the multipolar interactions up to the octupole terms are necessary to remove the macroscopic electric field effects in elastic constants at $\mathcal{O}(q^{2})$, while the higher-order terms such as the dipole-hexadecapole and quadrupole-octupole interactions are required to converge the bending rigidity tensor. For this purpose, we have further developed a reliable cluster expansion model to efficiently extract the high-order multipole and dielectric tensors. We then validate the effectiveness and accuracy of our methodology by demonstrating the example calculations on four bulk solids: Si, NaCl, GaAs and BaTiO3, and four 2D materials: graphene, h-BN, MoS2 and InSe, where an overall good agreement has been achieved among other theoretical methods as well as the experimental measurements. Interestingly, we discover that the lattice-mediated contributions due to the relaxation of internal atomic positions under bending deformation play a crucial role in stabilizing bent 2D materials and flexural modes, which has not yet been reported. While the lattice relaxation effect tends to lower the bending rigidities of the single-layer 2D materials, it can stiffen the bending modulus of multilayer 2D materials. Our findings will deepen the understanding of bending deformation of nanoplates and membranes, complementing the microscopic theory of flexural nanomechanics.

### v2

@@ IV.6 Elastic tensor and bending rigidity in semiconducting monolayers
To validate the generality of our bending rigidity approach, we also tested semiconducting monolayers, where long-range fields couple with lattice vibrations. We start with h-BN but remark that the calculation of quadrupole tensor in DFPT is not implemented in Quantum ESPRESSO [111, 112] and the recent generalization of charge density response to 2D systems is limited to in-plane atomic displacements [129]. Therefore, we use the quadrupole tensors of h-BN, MoS2 and InSe computed by some of us [105] using the DFPT implementation in the abinit package [165, 104, 166, 167] and reported in Table S1 of the SM [118]. To guarantee result consistency, we therefore use the same computational parameters as Ref. [105]. As a side remark, we find that the use of 2D Fourier interpolation with the long-range presented in Sec. III.4 combined with the rotational invariance and equilibrium conditions [97] gives small imaginary or linear ZA modes in the long-wavelength limit, due to out-of-plane dynamical dipoles, i.e. the second term in square parentheses in Eq. (79). We find that these out-of-plane dipoles have negligible contributions to both the phonon dispersions and elastic properties in the studied materials and leave this limitation for further studies. We give additional details on the rotational invariance and equilibrium conditions in Appendix A.
@@
Figure 6: (a) Phonon dispersion and (b) independent components of the elastic and bending rigidity tensors in h-BN as a function of supercell size. The multipolar interactions up to the quadrupole terms are considered for the long-range effect in the phonon dispersion of h-BN, and the inset of panel (a) shows the enlarged acoustic branches around the $\Gamma$ point along the K and M directions. In the panel (b), the short-circuit elastic constants and bending rigidities of h-BN are calculated at the levels of removing DD and DD+DQ+QQ interactions; the dotted lines represent the theoretical elastic constants $C_{11}$$C_{12}$ and bending rigidity $D_{11}$ taken from Ref. [157] and Ref. [90], respectively.
@@
Figure 7: Phonon dispersions and independent components of the elastic and bending rigidity tensors in InSe as a function of supercell size. The phonon spectra in the panels (a) and (b) are obtained based on a Fourier interpolation with the DD and DD+DQ+QQ interactions taken into account, respectively, and their insets show the enlarged acoustic branches around the $\Gamma$ point along the K and M directions. In the panel (c), the short-circuit elastic constants and bending rigidities of InSe are also calculated at the levels of removing DD and DD+DQ+QQ interactions; the dotted lines represent the theoretical elastic constants $C_{11}$ and $C_{12}$ taken from Ref. [164].
@@
When it comes to the bending rigidities of h-BN shown in Fig. 6(b), the convergence is faster than for elastic constants and reached for a $\times6$$\times$1 supercell, which is connected to fast convergence of the long-wavelength dispersion of ZA modes as shown in the inset of Fig. 6(a). However, we notice a small drift at higher supercell sizes for the bending rigidities because in monolayers the necessary condition of an $n$-order moment of IFCs to converge is that the IFCs themselves decay faster than $1/d^{n+2}$, as the integral in Eq. (73) is performed over the in-plane area of a supercell. This implies that the decay of short-range IFCs in real space must be faster than $1/d^{6}$ in order to converge the fourth-order moment of IFCs in 2D materials, which is equivalent to the requirement that all the multipolar interactions at $\mathcal{O}(q^{4})$ should be be removed; these include the first-order DD, the second-order DQ, the third-order QQ and DO, and the fourth-order QO and DH interactions, where H denotes the hexadecapoles. We leave the calculations of multipoles moments higher than quadrupoles in 2D for further studies but notice that our results range within 1.01–1.06 eV, which are still in good agreement with the existing theoretical results, 1.02 [90] and 0.95 eV [158]. Also, interestingly, the effects of DQ and QQ interactions on bending rigidities are negligible, as shown in Fig. 6(b). Besides, we predict for the first time $D_{12}=0.28$ eV and a Gaussian modulus of -0.79 eV in h-BN. Lastly, unlike graphene, the lattice relaxation does contribute to the bending rigidities of h-BN, that is, -0.11 eV for both $D_{11}$ and $D_{12}$, which is negative and thus lowers the energy of the system upon bending.
@@
We continue to study the case of 2D MoS2 monolayer, whose detailed results are shown in Fig. S4 of the SM [118]. We find that convergence for that material is easier because the macroscopic long-range effects are weak, which we attribute to the relatively large in-plane dielectric permittivity but accompanied with small effective charges. Based on the IFCs of a $\times12$$\times$1 supercell, we obtain the $C_{11}$ and $C_{12}$ of monolayer MoS2 as 131.88 and 32.28 N/m, respectively, while the previous DFT values are 132.3 and 32.8 N/m [161]. The detailed mechanical and elastic properties of 2D MoS2 are summarized in Table 4, in excellent agreement with experiment. In addition, the converged principal bending rigidity $D_{11}$ based on a $\times12$$\times$1 supercell is 9.06 eV, consistent with a value of 9.01 [89] and 9.10 eV [87], which both relies on the Helfrich Hamiltonian method. Notably, we find that the lattice-mediated contribution due to relaxation effects significantly impacts the bending rigidities of MoS2, which is a crucial difference from the behavior observed in graphene and h-BN. Without the lattice-mediated contribution, both $D_{11}$ and $D_{12}$ become negative (i.e. -1.53 and -0.51 eV), which indicates that such a material is not stable under a bending perturbation due to the failure of the clamped-ion approximation. It further suggests that the motions of bending waves in MoS2 can be only stabilized by the renormalization effect of the lattice, a phenomenon that has not been discussed before. Finally, we here report for the first time the values of $D_{12}$ and Gaussian modulus $D^{\mathrm{G}}$ of MoS2 are 4.35 and -4.71 eV, respectively, awaiting the future experimental verifications.
@@
Our last example for demonstrating the accuracy of the proposed method is monolayer InSe, which serves as a prototype for the 2D family of III–VI metal chalcogenides. The calculated phonon dispersions with the increasing supercell size are shown in Figs. 7(a) and (b), where the long-range Coulomb electrostatics is treated at the levels of DD and DD+DQ+QQ interactions, respectively. In the case of DD interaction only, the slope of the linear TA modes around the $\Gamma$ point changes slowly as the supercell size increases. As can be seen from Fig. 7(b) and the inset therein, when the DD+DQ+QQ interpolation scheme is used, the resulting phonon spectra can be readily converged using only a small $\times6$$\times$1 supercell. In particular, the long-wavelength behaviors of all three acoustic modes remain constant with the increasing supercell size from $\times4$$\times$1 to $\times14$$\times$1. For the short-circuit elastic constants, as illustrated in Fig. 7(c), the values of $C_{11}$ and $C_{12}$ are 49.60 and 20.43 N/m, respectively. While $C_{11}$ agrees well with the reference result of 49.38 N/m from the previous DFT calculations [164], $C_{12}$ shows a large error around 48.91% with respect to the reference value of 13.72 N/m. Such an inconsistency can be remedied by further taking into account the contributions from DQ and QQ interactions, as shown by the red dotted lines in Fig. 7(c). When the multipolar interactions up to the quadrupole terms are considered, values of 49.27 and 13.38 N/m are obtained, respectively, in excellent agreement with the previous reported values. Also, if the clamped-ion approximation is adopted, the respective $C_{11}$ and $C_{12}$ are 73.38 and 23.39 N/m. The detailed results of elastic constants and moduli for 2D InSe are summarized in Table 4.
@@
Similarly to the cases of h-BN and MoS2, the bending rigidities of InSe in Fig. 7(c) suffer from the issue of divergence when using a large supercell, due to the existence of high-order mutipolar interactions beyond the quadrupole terms. Although the removal of DQ and QQ interactions only results in a small positive shift in the $D_{12}$ component compared to the case with only DD interactions subtracted, the bending rigidities are still diverging, especially from the supercell size of $\times12$$\times$1. We note that the impact of DD term itself is also relatively small, see Fig. S5 of the SM [118], which suggests that the octupole tensor will be important in this material. However, owing to the technical difficulty of computing higher-order multipolar interactions, we will leave this point for the future investigation. Using a moderate $\times10$ supercell, our approximated principal bending rigidity $D_{11}$, Gaussian modulus $D^{\mathrm{G}}$, and the $D_{12}$ component of InSe monolayer are 22.60, -14.25, and 8.34 eV, respectively. By comparison, the principal bending rigidities of InSe from the Helfrich Hamiltonian method within the cyclic DFT calculations are 15.3 and 17.5 eV, respectively, using the roll-up nanotubes along the zigzag and armchair directions [89]. Such discrepancy could be attributed to many factors, including unconverged DFT parameters and numerical instability in extracting bending rigidities from the energy-curvature relation. The latter is a drawback of the Helfrich Hamiltonian method, which can be corroborated by the observed symmetry breaking in the calculated bending rigidities along the two principal axes. Furthermore, our simulation also reveals a significant renormalization effect from the lattice relaxation on bending rigidities of InSe, where the clamped-ion $D_{11}$ and $D_{12}$ are found to be 0.32 and 0.11 eV. Interestingly, for 2D materials containing only a single layer such as h-BN the lattice relaxation induced by a bending deformation tends to lower the bending rigidity, while it can significantly stiffen the bending modulus of multilayer 2D materials. This finding will provide practical guidelines for designing future flexible electronics.
@@ V Conclusions
In this work, we propose a novel method to simultaneously compute the elastic constants and bending rigidities of bulk and LD materials. Compared to existing methods, the approach is based on the long-wavelength perturbation theory and uses only the real-space IFCs, dynamical charges and dielectric tensors as input. This provides the direct benefit to work for any dimensions unlike the finite-difference method based on the strain-stress and strain-energy relations where the elastic constants involving the out-of-plane Cartesian component should be removed in the constitutive equations for 2D materials. Second, the electromechanical coupling is treated on the same footing through the calculated multipole and dielectric tensors. One can readily obtain the open-circuit elastic constants via the piezoelectric tensor calculated directly from the quadrupole tensor, with the lattice-mediated contribution fully included. Third, the bending rigidities of 2D materials can be easily obtained based on a higher-order long-wavelength perturbation expansion in the same manner. Notably, the calculated bending rigidity is inherently tensorial, which further allows us to explore the elastic anisotropy of 2D plates, and currently this is the only way to obtain the complete fourth-rank bending rigidity tensor, to the best of our knowledge. Fourth, temperature-dependent elastic constants and bending rigidities beyond the quasiharmonic approximation can be readily achieved via the self-consistent harmonic approximation [94] by taking into account the full quantum anharmonicity, as our approach takes the real-space IFCs as the input. Last, if phonon calculations have been already performed, the calculations of elastic constants and bending rigidities with the present approach can be performed at virtually no additional computational cost. This enables the automated high-throughput calculations of elastic properties directly through existing phonon workflows that have been already implemented in most of high-throughput workflow managers [168, 169, 170].
@@
In summary, we present an efficient and accurate approach for calculating both the elastic and bending rigidity tensors of crystalline solids, which is rigorously formulated based on the secular equation of lattice dynamics and the long-wavelength perturbation theory. In particular, we have provided a practical recipe in the framework of modern electronic structure calculations for implementing the Huang formula of elastic stiffness tensor that dates back to the atomic theory of elasticity introduced in 1950. Besides, we extend the Huang’s theory to describe also the equations of motion of bending wave in 2D materials and obtain the corresponding bending rigidity tensor, through the fourth-order long-wavelength expansion. Significantly, in semiconductors and insulators, the induced macroscopic electric fields couple with the lattice vibrations in the long-wavelength limit, which must be treated separately in order to recover the elastic constants and bending rigidities at the correct electrical boundary conditions. To obtain the short-circuit one, it is found that the multipolar interactions up to the octupole terms are necessary to remove the macroscopic electric field effects in elastic constants at $\mathcal{O}(q^{2})$, while the higher-order terms such as the dipole-hexadecapole and quadrupole-octupole interactions are required to converge the bending rigidity tensor. For this purpose, we have further developed a reliable cluster expansion model to efficiently extract the high-order multipole and dielectric tensors. We then validate the effectiveness and accuracy of our methodology by demonstrating the example calculations on four bulk solids: Si, NaCl, GaAs and BaTiO3, and four 2D materials: graphene, h-BN, MoS2 and InSe, where an overall good agreement has been achieved among other theoretical methods as well as the experimental measurements. Interestingly, we discover that the lattice-mediated contributions due to the relaxation of internal atomic positions under bending deformation play a crucial role in stabilizing bent 2D materials and flexural modes, which has not yet been reported. While the lattice relaxation effect tends to lower the bending rigidities of the single-layer 2D materials, it can stiffen the bending modulus of multilayer 2D materials. Our findings will deepen the understanding of bending deformation of nanoplates and membranes, complementing the microscopic theory of flexural nanomechanics.
Last but not least, the computational framework presented here serves as a robust foundation for identifying and characterizing promising energy materials. For instance, the accurate determination of mechanical descriptors, combined with high-throughput phonon calculations, can enable the efficient screening of high thermal conductivity materials or the rational design of strain-tolerant cathode materials for long-lifespan batteries. Furthermore, our automated extraction of high-order multipoles (particularly quadrupoles and octupoles) provides an atom-resolved pathway to optimize piezoelectric and flexoelectric materials. By bridging the gap between fundamental material properties and device-level engineering, these developments are crucial for the practical realization of next-generation energy harvesting and electromechanical devices.